OpenStudy (anonymous):

Add and Simplify: 4/m+2 + 2/m-2

OpenStudy (anonymous):

\[\frac{ 4 }{ m+2 } + \frac{ 2 }{ m-2 }\] =\[\frac{ 4 }{ m+2 } (\frac{ m-2 }{ m-2}) + \frac{ 2 }{ m-2 }(\frac{ m+2 }{ m+2 })\] =\[\frac{ 4m-8 }{ (m+2)(m-2) } + \frac{ 2m+4 }{ (m-2)(m+2) }\] =\[\frac{ 6m-4 }{ (m+2)(m-2)}\] or \[\frac{ 6m-4 }{ m^2-4}\] i think...